# Whispering in the dark

### v2

@@ 2.2 Particle acceleration
In collisionless environments, particles are susceptible to being accelerated up to relativistic speeds by shocks, turbulence, and magnetic reconnection (Gruzinov 1998).
@@ 2.2.1 Shocks and turbulence
The surroundings of an accreting black hole are prone to magnetised relativistic shocks due to the accretion-ejection dynamics. They are expected in black hole’s jets (Malzac 2013) and collimated disc outflows (Jacquemin-Ide et al. 2021). At shocks, particles can bounce back and forth across the shock and gain each time a certain amount of kinetic energy (see review by Pelletier et al. 2017). This process called diffusive shock acceleration (i.e. first-order Fermi process), can accelerate particles up to relativistic speeds (Ellison et al. 1990). Furthermore, black holes’ corona may show strong Alfvénic turbulence (Sandoval et al. 2023; Grošelj et al. 2024) where particles undergo stochastic acceleration (i.e. second-order Fermi process) through multiple magnetic mirroring episodes which provide them with a net amount of kinetic energy (Fermi 1949). Both processes accelerate ions and electrons into a non-thermal power-law energy distribution (Comisso & Sironi 2022). It has been shown to be a process susceptible to contributing to the heating of black holes’ coronae (Chandran et al. 2018; Hankla et al. 2022).
@@ 2.2.2 Magnetic reconnection
Particles can also be accelerated via magnetic reconnection which occurs in current sheets formed at the interface between magnetic field lines of opposite polarity. In these sheets, a very high non-ideal electric field appears at the X-points when magnetic field lines reconnect (Asenjo & Comisso 2019). It accelerates particles which gather into magnetic islands, called plasmoids, formed in the current sheet by the tearing instability (Zenitani & Hoshino 2007). The rate at which electromagnetic energy is dissipated and converted into particle kinetic energy (the reconnection rate) depends on the cold magnetisation parameter of the plasma, defined as the ratio of the magnetic to the inertial mass energy of the particles (for a recent review, see Kagan et al. 2015):
with the magnitude of the magnetic field, the plasma number density, the bulk Lorentz factor and the mass of the electron. Estimates for stellar-mass black holes indicate that the magnetic field in which the electrons are embedded can be strong enough to accelerate them up to relativistic Lorentz factors via magnetic reconnection in the corona (Poutanen & Vurm 2009). The efficiency of this mechanism is hardly affected by the presence or not of a disc, as shown by the spherically symmetric simulations ran by (Galishnikova et al. 2023) where reconnecting current sheets spontaneously form as the highly magnetised plasma is accreted. Furthermore, the reconnection rate is enhanced by turbulence (Lazarian & Vishniac 1999; Lazarian et al. 2012) and turbulence-driven magnetic reconnection has been shown to be an efficient acceleration mechanism at the basis of the black hole’s jet (Singh et al. 2015).
@@ 2.2.3 Relative contribution
The relative contribution of the above mechanisms to particle acceleration depends on the magnetisation parameter . For instance, particle acceleration at relativistic shocks is significantly quenched when is above (Lemoine & Pelletier 2010; Sironi et al. 2013; Plotnikov et al. 2018), while the magnetic reconnection rate plateaus at 10% when is above a few 10 (Werner et al. 2018).
We now evaluate the magnetisation parameter in the BH vicinity assuming that the stellar magnetic flux is advected inward by the accreted material. We work in the wind accretion regime, where the wind speed the orbital speed ( is the orbital angular speed) and the accretion radius the orbital separation. At the bow shock, formed upstream of the black hole by the gravitational beaming of the stellar wind, the density can be approximated by:
@@
where we introduced the parameter to represent the decrease of the accretion rate as we get closer to the BH due to outflows (Xie & Yuan 2012). Hereafter, in Sect. 3.2, we introduce the accretion radius which stands for the gravitational cross-section of the BH. As the stellar wind is beamed toward the BH, it forms a bow shock of size comparable to the accretion radius. Within this region, we assume that the plasma flowing onto the BH is spherically symmetric. Then, using Eq. (3) and the expression the accretion radius in Eq. (19), we deduce the magnetisation parameter of the electrons at , where is the gravitational radius from the BH, to be:
where we assumed that since the ions’ mass dominates and = 1. We consider a fiducial value of to account for the outflows within the shocked region around the BH (Xie & Yuan 2012). In these conditions, lower limits on can be obtained by setting to the terminal wind speed of an O star (typically 2,000 km s-1). We take the following standard values: a stellar mass loss rate typical of an LMC O-type star Myr-1 (Brands et al. 2022), a stellar magnetic field of G (30 G is below the detection sensitivity of 80% of the stars observed in the MiMeS survey, Wade et al. 2016; Grunhut et al. 2017; Petit et al. 2019), and a stellar radius of R⊙. Then, for orbital separations ranging between 3 R∗ and 100 R∗, is at least . The magnetisation parameter can also be estimated from observations interpreted via radiative models. For instance, in the BH-hosting HMXB Cygnus X-1, the analysis of the polarised synchrotron emission (Cangemi et al. 2021) indicates that in the innermost regions, which is coherent with the values obtained for a magnetic field near equipartition with the ion energy density (Malzac & Belmont 2009). This agrees with the values we derived from our toy model of stellar magnetic flux compression.
Given the high magnetisation parameters we can expect around stellar-mass BHs fed by the wind from an OB stellar companion, it is safe to assume that the main particle acceleration mechanism in the corona is magnetic reconnection in the relativistic regime (i.e. higher than a few 10), where the reconnection rate is 10% (or even higher if reconnection is turbulence-driven). In the MAD regime, the approximate equipartition between accretion and magnetic energy density implies that a large fraction of could serve to accelerate the electrons via magnetic reconnection. This is in agreement with observations of supermassive BHs where fits from detailed radiative models indicate that the electron heating parameter , that is, the fraction of the viscously dissipated energy that heats the electrons, is high, with values ranging between and (Yuan et al. 2003; Yu et al. 2011; Liu & Wu 2013).
@@ 2.3 Non-thermalised emission
In classic radiatively inefficient accretion flows, be it because of the lack of accretion disc or of the low , the ions and the electrons are no longer coupled through Coulomb interactions due to the low plasma density (Mościbrodzka et al. 2016). The temperatures of the ions and electrons differ and the electron distribution can develop a significant non-thermal component through particle acceleration. Indeed, Coulomb collisions between electrons become negligible for (Sharma et al. 2007), and therefore the presence of a non-thermal population is needed to account for the spectrum of Cygnus X-1 in the low/hard state for instance (Cangemi et al. 2021). We show in Sect. 2.2 that the energy reservoir contained in the magnetised accretion flow can be efficiently tapped through magnetic reconnection in the very sub-Eddington regime (see also (Sharma et al. 2007)). The electrons accelerated through this process will emit non-thermal radiation through synchrotron, bremsstrahlung, and inverse Compton scattering. Hereafter, we focus on the collisionless regime and estimate the synchrotron luminosity from the purely non-thermal population of electrons which should dominate at low mass accretion rate.
@@ 2.3.1 Synchrotron
where we used the ambient magnetic field and electron number density values in the denominator found by Cangemi et al. (2021) in the vicinity of the 20M⊙ BH in Cygnus X-1 (Miller-Jones et al. 2021), from the analysis of the polarised emission.
This idealised one-zone model does not account for the non-uniform density and magnetic field, nor the shape of the underlying distribution of electron energy. In the relativistic regime, the maximum Lorentz factor reachable by an electron accelerated by magnetic reconnection over a length scale in an ambient magnetic field is:
where is the plasma multiplicity. Even at low mass accretion rates and for a maximally rotating BH (), we expect (El Mellah et al. 2022), so in the magnetised plasma surrounding stellar-mass BHs capturing the wind from an OB companion.
Acceleration through magnetic reconnection is typically described by a power law distribution with an exponent (Werner et al. 2016; El Mellah et al. 2022). Hence, the contribution of the highest energy electrons dominate the emission. However, this simplified estimate of the synchrotron power with shows that, provided electrons are accelerated to significant Lorentz factors and embedded in a high magnetic field, they can produce a significant amount of non-thermal synchrotron emission. For stellar-mass BHs without an accretion disc, the peak of this emission is expected at a photon energy of 0.2 keV (with the values above of Lorentz factor and magnetic field), in soft X-rays.
@@ 2.3.2 Bremsstrahlung
For a stellar-mass BH accreting mass from the interstellar medium, it has been shown that the resulting luminosity is extremely low, orders of magnitude below observable limits (see discussion in Sect. 7.8, following Eq. (7.16) of Frank et al. 2002). For BHs accreting material from the wind of an OB star companion, the mass accretion rate is largely set by the wind mass loss rate of the OB star and the binary orbital period. The radiative luminosity in the X-ray band is at most erg s-1 at orbital period of 10 d and falls steeply with the mass accretion rate, such that at orbital periods of 100 d the luminosity is less than erg s-1 (Quast et al. in prep). Hence, thermal bremsstrahlung is not expected to contribute to observable X-ray emission.
X-ray line emissivity due to bound-bound transitions can be significantly higher than the thermal bremsstrahlung emission if the temperature of ions drops below 106K. However, this temperature regime should not be relevant near the BH since the ion temperature is much higher than the electron temperature in the magnetised plasma, and electron temperature is typically above 106K.
@@ 2.3.3 Inverse Compton scattering
Relativistic electrons upscatter soft X-ray photons through Inverse Compton in optically thin and hot accretion flow. The seed photons can come from the multi-colour black body emission from an underlying thin disc (at high mass and angular momentum accretion rates), or the synchrotron emission. In the latter case, it could shift the peak of synchrotron emission up to hard X-rays (Sridhar et al. 2021), although Comptonisation is expected to be subdominant when (Esin et al. 1997).
3 Methods
We discuss the grid of binary evolution models utilised for this work in Sect. 3.1. We discuss the criterion for forming an accretion disc around the BH in BH+OB star binaries in Sect. 3.2. We outline the procedure to calculate the X-ray luminosity from a BH+OB star binary in the presence and absence of a geometrically thin, optically thick accretion disc in Sect. 3.3 and Sect. 3.4, respectively. In Sect. 3.5, we define the distribution functions that are used to predict the observable stellar parameters from our binary models during the BH+OB star phase.
@@ 3.1 Stellar evolution models
Upon core carbon depletion of the primary, we assume that if the helium core mass of the primary is larger than 6.6 , the helium core of the primary directly collapses into a BH without a natal kick and the mass of the BH formed equals the helium core mass of the primary (Ertl et al. 2016; Müller et al. 2016; Sukhbold et al. 2018; Langer et al. 2020). The BH kick depends on the neutrino energy available and the asymmetry of the fall-back material (Belczynski et al. 2012; Janka 2013). Empirical evidence towards the magnitude of BH kicks remains inconclusive (e.g. Farr et al. 2011), with some studies positing the need for a low kick (Wong et al. 2012), while others requiring a significantly high kick (Repetto et al. 2012). It has been recently shown that a BH kick is not necessary to explain the lack of observed wind-fed BH HMXBs in the Milky Way compared to the number of Wolf-Rayet+O star binaries (Sen et al. 2021). When the first BH forms, the secondary (hereafter called the OB star companion to the BH) still burns hydrogen in its core. This marks the onset of the BH+OB star phase studied in this work.
@@
As the models undergo mass transfer before the BH+OB star phase and there is no kick velocity, we assume the orbit during the BH+OB star phase will remain circular. Lastly, we assume that the BHs formed from the collapse of the primary have negligible spins (Kerr parameter = 0, Qin et al. 2018; Ma & Fuller 2023). The BH+OB star phase ends when the OB star companion completes core hydrogen burning or fills its Roche lobe while on the main sequence.
3.2 Disc-formation criterion
To determine whether a disc forms, the specific angular momentum of the infalling matter onto the BH must be compared to the specific angular momentum of a test particle at the ISCO of the BH. For a spinning black hole, it is entirely set by the black hole mass and its dimensionless spin , with for a non-spinning black hole and for a maximally spinning black hole:
@@ 3.2 Disc-formation criterion
If accretion proceeds through stellar wind capture (instead of Roche lobe overflow), a bow shock forms around the black hole (El Mellah & Casse 2015). Within the shocked region, matter flattens into a disc, that is, a centrifugally maintained structure, provided its specific angular momentum is higher than . When the BH accretes matter from the OB star wind, and the wind speed at the orbital separation is larger than the orbital speed , we have (c.f. Eq. (10) of Sen et al. 2021, where the criterion is expressed in terms of the ratio of the circularisation radius of a Keplerian accretion disc to the innermost stable circular orbit of the BH)
@@
Equation (18) shows that accretion discs form preferentially in systems where the wind speed is low compared to the orbital speed. We can express the condition for disc formation as
where is the orbital period. Hydrodynamical simulations of accretion onto stellar-mass BHs suggest that (Livio et al. 1986; Ruffert 1999). For a Schwarzschild BH, (Carroll 2019). In the Newtonian approximation, for a non-spinning BH. In Sect. 4, we show the results for to be consistent with the accretion disc formation criterion in Sen et al. (2021), but also discuss the statistics for in Sect. 6.4.
3.3 X-ray luminosity from a Keplerian accretion disc
If an accretion disc can form, the X-ray luminosity from the accretion disc is calculated as

### v3

@@ 2.2 Particle acceleration
In collisionless environments, particles are susceptible to being accelerated up to relativistic speeds by shocks, turbulence and magnetic reconnection (Gruzinov 1998).
@@ 2.2.1 Shocks and turbulence
The surroundings of an accreting black hole are prone to magnetized relativistic shocks due to the accretion-ejection dynamics. They are expected in black hole’s jets (Malzac 2013) and collimated disk outflows (Jacquemin-Ide et al. 2021). At shocks, particles can bounce back and forth across the shock and gain each time a certain amount of kinetic energy (see review by Pelletier et al. 2017). This process called diffusive shock acceleration (i.e. first-order Fermi process), can accelerate particles up to relativistic speeds (Ellison et al. 1990). Furthermore, black holes’ corona may show strong Alfvénic turbulence (Sandoval et al. 2023; Grošelj et al. 2024) where particles undergo stochastic acceleration (i.e., second-order Fermi process) through multiple magnetic mirroring episodes which provide them with a net amount of kinetic energy (Fermi 1949). Both processes accelerate ions and electrons into a non-thermal power-law energy distribution (Comisso & Sironi 2022). It has been shown to be a process susceptible to contributing to the heating of black holes’ coronae (Chandran et al. 2018; Hankla et al. 2022).
@@ 2.2.2 Magnetic reconnection
Particles can also be accelerated via magnetic reconnection which occurs in current sheets formed at the interface between magnetic field lines of opposite polarity. In these sheets, a very high non-ideal electric field appears at the X-points when magnetic field lines reconnect (Asenjo & Comisso 2019). It accelerates particles which gather into magnetic islands, called plasmoids, formed in the current sheet by the tearing instability (Zenitani & Hoshino 2007). The rate at which electromagnetic energy is dissipated and converted into particle kinetic energy (the reconnection rate) depends on the cold magnetization parameter of the plasma, defined as the ratio of the magnetic to the inertial mass energy of the particles (for a recent review, see Kagan et al. 2015):
with the magnitude of the magnetic field, the plasma number density, the bulk Lorentz factor and the mass of the electron. Estimates for stellar-mass black holes indicate that the magnetic field in which the electrons are embedded can be strong enough to accelerate them up to relativistic Lorentz factors via magnetic reconnection in the corona (Poutanen & Vurm 2009). The efficiency of this mechanism is hardly affected by the presence or not of a disk, as shown by the spherically-symmetric simulations ran by (Galishnikova et al. 2023) where reconnecting current sheets spontaneously form as the highly magnetized plasma is accreted. Furthermore, the reconnection rate is enhanced by turbulence (Lazarian & Vishniac 1999; Lazarian et al. 2012) and turbulence-driven magnetic reconnection has been shown to be an efficient acceleration mechanism at the basis of the black hole’s jet (Singh et al. 2015).
@@ 2.2.3 Relative contribution
The relative contribution of the above mechanisms to particle acceleration depends on the magnetization parameter . For instance, particle acceleration at relativistic shocks is significantly quenched when is above (Lemoine & Pelletier 2010; Sironi et al. 2013; Plotnikov et al. 2018), while the magnetic reconnection rate plateaus at 10% when is above a few 10 (Werner et al. 2018).
We now evaluate the magnetization parameter in the BH vicinity assuming that the stellar magnetic flux is advected inward by the accreted material. We work in the wind accretion regime, where the wind speed the orbital speed ( is the orbital angular speed) and the accretion radius the orbital separation. At the bow shock, formed upstream of the black hole by the gravitational beaming of the stellar wind, the density can be approximated by:
@@
where we introduced the parameter to represent the decrease of the accretion rate as we get closer to the BH due to outflows (Xie & Yuan 2012). Hereafter, in Sect. 3.2, we introduce the accretion radius which stands for the gravitational cross-section of the BH. As the stellar wind is beamed toward the BH, it forms a bow shock of size comparable to the accretion radius. Within this region, we assume that the plasma flowing onto the BH is spherically symmetric. Then, using equation (3) and the expression the accretion radius in equation (19), we deduce the magnetization parameter of the electrons at , where is the gravitational radius from the BH, to be:
where we assumed that since the ions’ mass dominates and = 1. We consider a fiducial value of to account for the outflows within the shocked region around the BH (Xie & Yuan 2012). In these conditions, lower limits on can be obtained by setting to the terminal wind speed of an O star (typically 2,000 km s-1). We take the following standard values: a stellar mass loss rate typical of an LMC O-type star Myr-1 (Brands et al. 2022), a stellar magnetic field of G (30 G is below the detection sensitivity of 80% of the stars observed in the MiMeS survey, Wade et al. 2016; Grunhut et al. 2017; Petit et al. 2019), and a stellar radius of R⊙. Then, for orbital separations ranging between 3 R∗ and 100 R∗, is at least . The magnetization parameter can also be estimated from observations interpreted via radiative models. For instance, in the BH-hosting HMXB Cygnus X-1, the analysis of the polarized synchrotron emission (Cangemi et al. 2021) indicates that in the innermost regions, which is coherent with the values obtained for a magnetic field near equipartition with the ion energy density (Malzac & Belmont 2009). This agrees with the values we derived from our toy model of stellar magnetic flux compression.
Given the high magnetization parameters we can expect around stellar-mass BHs fed by the wind from an OB stellar companion, it is safe to assume that the main particle acceleration mechanism in the corona is magnetic reconnection in the relativistic regime (i.e. higher than a few 10), where the reconnection rate is 10% (or even higher if reconnection is turbulence-driven). In the MAD regime, the approximate equipartition between accretion and magnetic energy density implies that a large fraction of could serve to accelerate the electrons via magnetic reconnection. This is in agreement with observations of supermassive BHs where fits from detailed radiative models indicate that the electron heating parameter , that is, the fraction of the viscously dissipated energy that heats the electrons, is high, with values ranging between and (Yuan et al. 2003; Yu et al. 2011; Liu & Wu 2013).
@@ 2.3 Non-thermalised emission
In classic radiatively-inefficient accretion flows, be it because of the lack of accretion disk or of the low , the ions and the electrons are no longer coupled through Coulomb interactions due to the low plasma density (Mościbrodzka et al. 2016). The temperatures of the ions and electrons differ and the electron distribution can develop a significant non-thermal component through particle acceleration. Indeed, Coulomb collisions between electrons become negligible for (Sharma et al. 2007), hence the presence of a non-thermal population to account for the spectrum of Cygnus X-1 in the low/hard state for instance (Cangemi et al. 2021). We have shown in Sect. 2.2 that the energy reservoir contained in the magnetized accretion flow can be efficiently tapped through magnetic reconnection in the very sub-Eddington regime (see also (Sharma et al. 2007)). The electrons accelerated through this process will emit non-thermal radiation through synchrotron, bremsstrahlung, and inverse Compton scattering. Hereafter, we focus on the collisionless regime and estimate the synchrotron luminosity from the purely non-thermal population of electrons which should dominate at low mass accretion rate.
@@ 2.3.1 Synchrotron
where we used the ambient magnetic field and electron number density values in the denominator found by Cangemi et al. (2021) in the vicinity of the 20M⊙ BH in Cygnus X-1 (Miller-Jones et al. 2021), from the analysis of the polarized emission.
This idealized one-zone model does not account for the non-uniform density and magnetic field, nor the shape of the underlying distribution of electron energy. In the relativistic regime, the maximum Lorentz factor reachable by an electron accelerated by magnetic reconnection over a length scale in an ambient magnetic field is:
where is the plasma multiplicity. Even at low mass accretion rates and for a maximally rotating BH (), we expect (El Mellah et al. 2022), so in the magnetized plasma surrounding stellar-mass BHs capturing the wind from an OB companion.
Acceleration through magnetic reconnection is typically described by a power law distribution with an exponent (Werner et al. 2016; El Mellah et al. 2022). Hence, the contribution of the highest energy electrons dominate the emission. However, this simplified estimate of the synchrotron power with shows that, provided electrons are accelerated to significant Lorentz factors and embedded in a high magnetic field, they can produce a significant amount of non-thermal synchrotron emission. For stellar-mass BHs without an accretion disk, the peak of this emission is expected at a photon energy of 0.2 keV (with the values above of Lorentz factor and magnetic field), in soft X-rays.
@@ 2.3.2 Bremsstrahlung
For a stellar-mass BH accreting mass from the interstellar medium, it has been shown that the resulting luminosity is extremely low, orders of magnitude below observable limits (see discussion in section. 7.8, following equation. (7.16) of Frank et al. 2002). For BHs accreting material from the wind of an OB star companion, the mass accretion rate is largely set by the wind mass loss rate of the OB star and the binary orbital period. The radiative luminosity in the X-ray band is at most erg s-1 at orbital period of 10 d and falls steeply with the mass accretion rate, such that at orbital periods of 100 d the luminosity is less than erg s-1 (Quast et al. in prep). Hence, thermal bremsstrahlung is not expected to contribute to observable X-ray emission.
X-ray line emissivity due to bound-bound transitions can be significantly higher than the thermal bremsstrahlung emission if the temperature of ions drops below 106K. However, this temperature regime should not be relevant near the BH since the ion temperature is much higher than the electron temperature in the magnetized plasma, and electron temperature is typically above 106K.
@@ 2.3.3 Inverse Compton scattering
Relativistic electrons upscatter soft X-ray photons through Inverse Compton in optically thin and hot accretion flow. The seed photons can come from the multi-colour black body emission from an underlying thin disk (at high mass and angular momentum accretion rates), or the synchrotron emission. In the latter case, it could shift the peak of synchrotron emission up to hard X-rays (Sridhar et al. 2021), although Comptonisation is expected to be sub-dominant when (Esin et al. 1997).
3 Method
We discuss the grid of binary evolution models utilised for this work in Sect. 3.1. We discuss the criterion for forming an accretion disk around the BH in BH+OB star binaries in Sect. 3.2. We outline the procedure to calculate the X-ray luminosity from a BH+OB star binary in the presence and absence of a geometrically-thin, optically-thick accretion disk in Sect. 3.3 and Sect. 3.4, respectively. In Sect. 3.5, we define the distribution functions that are used to predict the observable stellar parameters from our binary models during the BH+OB star phase.
@@ 3.1 Stellar evolution models
Upon core carbon depletion of the primary, we assume that if the helium core mass of the primary is larger than 6.6 , the helium core of the primary directly collapses into a BH without a natal kick and the mass of the BH formed equals the helium core mass of the primary (Ertl et al. 2016; Müller et al. 2016; Sukhbold et al. 2018; Langer et al. 2020). The BH kick depends on the neutrino energy available and the asymmetry of the fall-back material (Belczynski et al. 2012; Janka 2013). Empirical evidence towards the magnitude of BH kicks remains inconclusive (e.g., Farr et al. 2011), with some studies positing the need for a low kick (Wong et al. 2012), while others requiring a significantly high kick (Repetto et al. 2012). It has been recently shown that a BH kick is not necessary to explain the lack of observed wind-fed BH HMXBs in the Milky Way compared to the number of Wolf-Rayet+O star binaries (Sen et al. 2021). When the first BH forms, the secondary (hereafter called the OB star companion to the BH) still burns hydrogen in its core. This marks the onset of the BH+OB star phase studied in this work.
@@
Since the models undergo mass transfer before the BH+OB star phase and there is no kick velocity, we assume the orbit during the BH+OB star phase will remain circular. Lastly, we assume that the BHs formed from the collapse of the primary have negligible spins (Kerr parameter = 0, Qin et al. 2018; Ma & Fuller 2023). The BH+OB star phase ends when the OB star companion completes core hydrogen burning or fills its Roche lobe while on the main sequence.
3.2 Disk formation criterion
To determine whether a disk forms, the specific angular momentum of the infalling matter onto the BH must be compared to the specific angular momentum of a test particle at the ISCO of the BH. For a spinning black hole, it is entirely set by the black hole mass and its dimensionless spin , with for a non-spinning black hole and for a maximally spinning black hole:
@@ 3.2 Disk formation criterion
If accretion proceeds through stellar wind capture (instead of Roche lobe overflow), a bow shock forms around the black hole (El Mellah & Casse 2015). Within the shocked region, matter flattens into a disk, that is, a centrifugally-maintained structure, provided its specific angular momentum is higher than . When the BH accretes matter from the OB star wind, and the wind speed at the orbital separation is larger than the orbital speed , we have (c.f. equation (10) of Sen et al. 2021, where the criterion is expressed in terms of the ratio of the circularisation radius of a Keplerian accretion disk to the innermost stable circular orbit of the BH)
@@
Equation (18) shows that accretion disks form preferentially in systems where the wind speed is low compared to the orbital speed. We can express the condition for disk formation as
where is the orbital period. Hydrodynamical simulations of accretion onto stellar-mass BHs suggest that (Livio et al. 1986; Ruffert 1999). For a Schwarzschild BH, (Carroll 2019). In the Newtonian approximation, for a non-spinning BH. In Sect. 4, we show the results for to be consistent with the accretion disk formation criterion in Sen et al. (2021), but also discuss the statistics for in Sect. 6.4.
3.3 X-ray luminosity from a Keplerian accretion disk
If an accretion disk can form, the X-ray luminosity from the accretion disk is calculated as